A person received 4% nominal interest. The inflation rate was -2% and the tax rate was 25%. This person received an after-tax real interest rate of 5%.

Group of answer choices

True

False
Answer #1

True

Nominal interest rate = 4%=0.04

Tax rate = 25%

So, the after tax return = 0.04×(1-0.25) = 0.03 i.e. 3%

Real interest rate = (1+0.03)/(1-0.02) - 1

= 0.051

=5.1% ~5%

Hope this helps. Do hit the thumbs up. Cheers!